# Fino1: On the Transferability of Reasoning-Enhanced LLMs to Finance

### v2

@@ 2.1 Datasets
Dataset Size Data Types Average Token
FinQA 1100 Tables and Texts 1,128
DM-Simplong 100 Tables and Texts 4,330
DM-Complong 300 Tables and Texts 39,983
@@
XBRL-Math 90 Texts and Equations 397
Table 1: Overview of the datasets used in the study.
More details of the datasets are listed in Table 1. By conducting experiments on these three datasets, we aim to evaluate the model’s ability in various aspects, including understanding financial terms, extracting relevant numbers and entities from financial reports from different sources, and reasoning over long contexts and multiple tables.
@@ 2.2 Evaluated Models
[…]
The evaluated models are summarized as follows: GPT-4o Hurst et al. (2024): OpenAI’s flagship model with multimodal capabilities and improved real-time reasoning. GPT-o1 Jaech et al. (2024): A reasoning-optimized variant in the GPT series with strong step-by-step problem-solving ability. GPT-o3-mini: A compact version of GPT-o3, optimized for efficiency with robust language understanding. DeepSeek-V3 Liu et al. (2024): A 671B Mixture-of-Experts (MoE) model with 37B active parameters per token, designed for efficient reasoning. DeepSeek-R1 Guo et al. (2025): A first-generation reasoning model with competitive performance on math, code, and reasoning tasks. Qwen2.5-72B-Instruct Yang et al. (2024): A strong multilingual model with advanced instruction tuning. Qwen2.5-72B-Instruct-Math: A math-specialized variant fine-tuned for complex problem solving. LLaMA-series Dubey et al. (2024): Widely adopted open-source models (8B–70B) with strong instruction-tuned variants used in our experiments. DeepSeek-R1-Distill Guo et al. (2025): Distilled models from DeepSeek-R1, fine-tuned using LLaMA and Qwen backbones to balance performance and efficiency. Other reasoning-enhanced models: Includes S1-32BMuennighoff et al. (2025), Limo Ye et al. (2025), and FinR1-7B Liu et al. (2025b), incorporating reasoning strategies such as financial CoT construction and task-specific fine-tuning.
Model Name Parameters Enhanced for Reasoning Tasks Context Window Size Close/Open Source Reasoning Enhanced Training Data
GPT-4o Unknown No 128k Closed -
GPT-o1 Unknown Yes 128k Closed Public and proprietary data (Human-annotated CoT, MCTS-assisted Synthetic data)
GPT-o3-mini Unknown Yes 128k Closed Public and proprietary data
DeepSeek-V3 671B No 128k Open -
DeepSeek-R1 671B Yes 128k Open Cold-start data generation, post-processing data
GPT-4.5 Unknown No 128k Closed -
Qwen2.5-72B-Instruct 72B No 128k Open -
Qwen2.5-72B-Instruct-Math 72B Yes 128k Open Synthetic data from Qwen, high-quality mathematical data, CoT, TIR
DeepSeek-R1-Distill-Llama-70B 70B Yes 128k Open Distilled from R1
Llama3-70B-Instruct 70B No 8k Open -
Llama3.1-70B-Instruct 70B No 128k Open -
Llama3.3-70B-Instruct 70B No 128k Open -
Qwen2.5-32B-Instruct 32B No 128k Open -
DeepSeek-R1-Distill-Qwen-32B 32B Yes 128k Open Distilled from R1
Limo 32B Yes 128k Open Public and proprietary data with Qwen filteration
S1-32B 32B Yes 128k Open Public and proprietary data + original data, filtered by quality, difficulty, diversity
Qwen/QwQ-32B 32B Yes 128k Open -
Qwen2.5-14B-Instruct 14B No 128k Open -
DeepSeek-R1-Distill-Qwen-14B 14B Yes 128k Open Distilled from R1
DeepSeek-R1-Distill-Llama-8B 8B Yes 128k Open Distilled from R1
Llama3-8B-Instruct 8B No 8k Open -
Llama3.1-8B-Instruct 8B No 128k Open -
Qwen2.5-7B-Instruct 7B No 128k Open -
FinR1-7B 7B Yes 128k Open 60k CoT data for financial tasks
Table 2: Summary of evaluated large language models, including their parameter sizes, reasoning capabilities, input limits, source availability, and reasoning enhancement strategies. MCTS refers to Monte Carlo Tree Search. CoT refers to chain-of-thought reasoning. TIR means tool-integrated reasoning.
@@ 2.3 Evaluation Settings
In each dataset, LLMs are presented with a context containing financial reports, tables, equations, and symbols, along with a question. Following prior work Zhao et al. (2024a), we use a generic prompt (see appendix D for more details) to instruct models to answer based on the given context. Models must extract relevant information and perform mathematical reasoning using numbers and concepts from the context. While outputs are typically in mathematical form, variations in formats (e.g., percentages, rounding) may cause differences. Reasoning-focused models like DeepSeek-R1 and GPT-o1 often generate step-by-step reasoning. Inspired by DocMath Zhao et al. (2024a), we adopt an LLM-as-judge evaluation approach: GPT-3.5-Turbo extracts final answers, followed by a mathematical comparison to assess correctness.
For large models (e.g., DeepSeek-V3, DeepSeek-R1, GPT series), we run evaluations via APIs from TogetherAI [^2] https://www.together.ai and OpenAI [^3] https://platform.openai.com, using versions GPT-o1-2024-12-17, GPT-o3-mini-2025-01-31, and GPT-4o-2024-08-06. For models up to 72B parameters, we use the VLLM framework on 8 A100 GPUs (80GB each), with a max generation length of 512 tokens.
2.3.1 Reasoning Data Curation
To enhance domain-specific reasoning capabilities in LLMs, we curate a high-quality training dataset by integrating diverse financial reasoning benchmarks. Our dataset combines a wide range of reasoning tasks—covering numerical computation, table-text integration, cross-table correlation, and long-context understanding—reflecting the complexity of real-world financial documents.
Specifically, we incorporate data from several well-established benchmarks, including FinQA Chen et al. (2022), DocFinQA Reddy et al. (2024b), TAT-QA Zhu et al. (2021), DocMath-Eval Zhao et al. (2024b), and BizBench-QA Krumdick et al. (2024). These datasets collectively provide rich and diverse coverage across task types and difficulty levels, supporting the development of robust financial reasoning models. FinQA provides QA pairs combining textual and tabular financial data; DocFinQA extends this to long-context documents; TAT-QA focuses on complex numerical reasoning; DocMath-Eval (CompLong and SimpShort subsets) targets reasoning over multi-table and single-table scenarios; and BizBench-QA enriches the corpus with code-based financial reasoning tasks. Please see Appendix B for detailed descriptions of each dataset.
2.4 Data Collection and Reasoning Path Construction
To ensure a balanced dataset of both short and long context samples, we collect all data from the FinQA train set, TATQA, DocMath-CompLong, and DocMath-SimpShort. Additionally, we sample 1,000 examples each from BizBench-QA and DocFinQA to balance short and long reasoning tasks. In total, we gather 9,851 financial QA samples for reasoning path construction. Inspired by Ye et al. (2025); Muennighoff et al. (2025), which suggest that simple QA pairs may not effectively improve reasoning capabilities, we apply a filtering strategy to retain more challenging examples. Specifically, we use Qwen-14B-Instruct to generate answers for each question and employ GPT-3.5-Turbo as a verifier. QA pairs correctly answered by Qwen-14B are removed, resulting in a more difficult training subset of 3,472 QA pairs for optimizing more advanced models.
2.5 Reasoning Path Generation and Refinement
For reasoning path generation, we use GPT-4o, chosen for its strong performance across datasets (Table 3) and its stability in generating high-quality outputs. Following the CoT refinement framework from HuatuoGPT-o1 Chen et al. (2024), we employ a verifier-based iterative process to refine the reasoning paths generated by GPT-4o. If a reasoning path is rejected by the verifier, refinement is guided by one of the following strategies: (1) Backtracking: Revisiting previous steps to correct logical inconsistencies. (2) Exploring New Paths: Generating alternative reasoning routes when the current path fails. (3) Verification: Cross-checking intermediate steps to ensure logical soundness. (4) Correction: Making targeted adjustments to rectify specific errors. The CoT refinement process begins with an initial prompt to generate a reasoning path and final answer. If the verifier identifies the answer as incorrect, one of the above strategies is applied iteratively until a correct and logically consistent answer is produced.
2.5.1 Model Training
We develop our financial reasoning models based on curated reasoning paths, focusing on smaller-sized models for efficient real-world deployment. We select LLaMA-3.1-8B-Instruct and Qwen2.5-14B-Instruct as backbone models, given their strong performance among small and mid-sized LLMs. We adopt a two-stage training approach. In the first stage, we perform Supervised Fine-Tuning (SFT), using questions as input and reasoning paths with final answers as output. For the 8B model, we use only FinQA reasoning paths to assess the impact of domain-specific fine-tuning. For the 14B model, we use the full set of sampled and filtered data to leverage its greater capacity for diverse and complex reasoning. In the second stage, we apply RL using a verifier-based sparse reward mechanism and optimize with Proximal Policy Optimization (PPO), following Chen et al. (2024). A reward model evaluates the correctness and coherence of the reasoning path. If errors are found, the model iteratively refines its outputs through backtracking, alternative path exploration, verification, and correction, improving logical consistency and equation-solving accuracy. During SFT, models are trained for three epochs with a learning rate of 5e-6 and a maximum sequence length of 8192 tokens. LoRA is applied for efficient parameter updates. In the RL stage, training continues for three epochs with a reduced learning rate of 5e-7, guided by the reward model to refine reasoning quality via PPO.
@@ 3 Results
Table 3 presents the performance of various LLMs across four financial reasoning benchmarks: FinQA, DM-Simplong, XBRL-Math, and DM-Complong. Our models, Fino1-8B and Fino1-14B, demonstrate strong and consistent results across tasks, outperforming several larger and closed-source models. Fino1-14B achieves an average score of 60.25, surpassing many larger-scale models such as LLaMA-3.3-70B-Instruct (56.04), Qwen2.5-72B-Instruct (53.71), GPT-o1-preview (54.05), GPT-o3-mini (57.89), and even GPT-4.5 (60.43). It also achieves the best performance among all models on two benchmarks—DM-Simplong (60.00) and XBRL-Math (86.67). These results demonstrate that high-quality, domain-specific reasoning training can enable smaller models to rival—and even surpass—the performance of significantly larger and proprietary LLMs.
Fino1-8B also performs competitively with an average score of 50.77, surpassing other 8B models and even some 14B and 32B models. Both models excel on XBRL-Math, where Fino1-14B achieves the top score of 86.67. Moreover, Fino1-14B and Fino1-8B significantly outperform FinR1-7B, a financial reasoning-enhanced model trained on a much larger 60K task-aligned dataset, demonstrating that larger data volume does not necessarily lead to better performance, and that carefully curated, high-quality reasoning data—even in smaller quantities—can yield superior results.
@@
Meta-Llama-3-70B-Instruct 58.92 41.00 56.67 13.67 42.57
@@
Fino1-8B 60.87 40.00 82.22 20.00 50.77
Fino1-14B 70.01 60.00 86.67 24.33 60.25
Table 3: Performance of different LLMs on four financial reasoning datasets.
Beyond overall performance, the comparison across models provides additional insights that can guide the development of LLMs for financial applications:
(1) General reasoning enhancements do not necessarily improve financial task performance. Models like GPT-o1 and DeepSeek-R1, though strong in general reasoning, underperform on financial benchmarks, highlighting the importance of domain-specific knowledge and structured numerical reasoning. Moreover, reasoning enhancement strategies based on general-domain data, as seen in S1-32B and Limo, often fail to improve and sometimes even degrade performance compared to their backbone models (e.g., Qwen2.5-32B). These results underscore that high-quality, domain-aligned reasoning paths are more effective than general reasoning augmentation, and tailored financial training is crucial for optimal model performance.
(2) Reasoning enhancements vary in effectiveness across financial tasks. Models like GPT-o1 perform worse than GPT-4o on FinQA (49.07 vs. 72.49) and DM-Simplong (56.00 vs. 60.00). Qwen2.5-Math improves in XBRL-Math (67.78 $\rightarrow$ 83.33) but drops sharply in FinQA (69.74 $\rightarrow$ 42.00) and DM-Simplong (59.00 $\rightarrow$ 42.00). DeepSeek-R1-Distill-Llama-8B also excels in XBRL-Math (81.11) but shows limited gains elsewhere. These results suggest that strategies like CoT or math fine-tuning may offer benefits for structured tasks but often fail to generalize effectively to broader financial reasoning. Compared with Qwen2.5-72B-Instruct and its math enhanced version, we can see that while TIR and synthetic building appears to help in equation-heavy tasks like XBRL-Math, its impact on other types of financial tasks is less clear and may be influenced by multiple factors such as task structure, data modality, or domain complexity.
(3) Larger model size offers limited benefits in financial reasoning. While larger models typically excel in general NLP tasks, their gains in financial tasks are not proportional. For example, DeepSeek-70B (66.73 FinQA, 53 DM-Simplong, 86.67 XBRL-Math) performs similarly to the smaller DeepSeek-R1-Distill-Qwen-32B (65.48, 55, 84.44), despite having less than half the parameters. Across models like Qwen2.5-72B, DeepSeek-R1-Distill-Llama-70B, and GPT-4o, performance tends to plateau around 70B. This may reflect the structured nature of tasks like XBRL-Math, where arithmetic precision outweighs complex language reasoning, enabling smaller models to compete effectively.
(4) Improved pre-training data and strategies in general domains also benefit financial tasks. Llama 3.1-8B-Instruct shows a notable performance boost over Llama 3-8B-Instruct (avg. 50.12 vs. 39.95), with gains in XBRL-Math (62.22 vs. 48.89) and FinQA (54.13 vs. 41.97). Similarly, Llama 3.1-70B-Instruct outperforms Llama 3-70B (58.17 vs. 52.20), especially in DM-Simplong (48.00 vs. 41.00), while Llama 3.3-70B-Instruct achieves the highest overall performance (64.05), driven by improvements in FinQA (68.15) and XBRL-Math (70.00). These gains stem from larger, more diverse training corpora (e.g., 15T+ tokens in Llama 3.1) and enhanced post-training techniques like online preference optimization. Such strategies improve general reasoning while also boosting performance in domain-specific tasks like financial reasoning.
(5) Existing reasoning models struggle with long contexts and cross-table correlations. Financial tasks often require integrating information from lengthy reports, structured tables, and equations, yet current models show clear limitations. For example, GPT-4o performs well on FinQA (72.49) but drops on DM-Simplong (60.00), indicating challenges in combining textual and symbolic reasoning. Models like DeepSeek-R1 and DeepSeek-R1-Distill-Qwen-32B excel in XBRL-Math but fail to generalize. Performance on DM-Complong remains low across models (e.g., LLaMA-3.3-70B: 32.00). Smaller models like DeepSeek-R1-Distill-Llama-8B also perform poorly with 33.00 for DM-Simplong. These results highlight the need for improved long-context modeling and structured data integration in financial reasoning.
(6) Financial reasoning enhancement improves performance across all tasks. Compared to LLaMA 3.1-8B-Instruct (41.16), DeepSeek-R1-Distill-Llama-8B shows a modest gain (43.94). In contrast, our Fino1-8B, trained on GPT-4o-distilled reasoning paths with backtracking strategies, achieves a substantial improvement, reaching an average score of 50.77, and consistently outperforms baselines on FinQA (60.87 vs. 54.13), DM-Simplong (40.00 vs. 34.00), and XBRL-Math (82.22 vs. 62.22). Fino1-14B further boosts performance, achieving an average score of 60.25, outperforming other 14B models (e.g., Qwen2.5-14B: 52.72) and even rivaling state-of-the-art models like DeepSeek-R1 (60.87) and GPT-4o (61.01). Notably, although Fino1-8B and Fino1-14B were trained with limited financial CoTs, both models show improvements across all tasks, highlighting the effectiveness of high-quality reasoning paths and structured training strategies such as backtracking, path exploration, and verification.
Overall, these insights highlight that financial reasoning poses unique challenges distinct from general NLP and mathematical reasoning. The performance variability across tasks underscores the need for specialized training strategies and architectural improvements to better support financial comprehension, structured data processing, and multi-step numerical reasoning.
@@ 4 Conclusion
This study presents a comprehensive evaluation of 24 general-purpose and reasoning-augmented LLMs on financial reasoning tasks, encompassing financial text understanding, numerical computation, and structured data interpretation. Our findings show that existing models, despite strong performance on general reasoning benchmarks, often fail to generalize to financial domains—particularly in handling long contexts, multi-table reasoning, and domain-specific terminology. To bridge this gap, we curated a high-quality, diverse financial reasoning dataset and introduced Fino1, a suite of domain-adapted models trained with GPT-4o-generated reasoning paths using chain-of-thought fine-tuning and reinforcement learning. Despite being trained on limited data, Fino1 models outperform several larger and advanced LLMs, including GPT-o1, GPT-o3-mini, GPT-4.5, and DeepSeek models (V3 and R1). These results highlight the effectiveness of domain-specific reasoning strategies and demonstrate the practical value of tailored model adaptation for real-world financial applications.
@@ Appendix A Related Work
A.1 General-Purpose and Reasoning-Enhanced LLMs
Recent years have witnessed significant advancements in large language models which demonstrated remarkable capabilities across various tasks, including natural language understanding, generation, and complex reasoning. Building upon these foundations, researchers have developed specialized reasoning-enhanced models such as DeepSeek-R1 and GPT-o1, which incorporate techniques like chain-of-thought prompting Wei et al. (2022), self-reflection Zhang et al. (2023), and reinforcement learning from human feedback Ouyang et al. (2022). These models have shown superior performance in mathematical reasoning Lewkowycz et al. (2022), symbolic manipulation Drori et al. (2023), and logical inference tasks Chen et al. (2023).

### v3

@@ 4.1 Tasks
Dataset Size Data Types Data Source Avg. Token Metric License
FinQA chen2021finqa 1,100 Tables and Texts Financial reports 1,128 Acc MIT License
DM-Simplong zhao2024docmath 100 Tables and Texts Financial reports 4,330 Acc MIT License
DM-Complong zhao2024docmath 300 Tables and Texts Financial reports 39,983 Acc MIT License
@@
XBRL-Math wang2025finnlp 90 Texts and Equations XBRL 397 Acc CDLA-Permissive 2.0
@@ 4.2 Evaluated Models
To investigate how existing LLMs, especially reasoning-enhanced LLMs, perform in financial tasks, we select a diverse set of 29 LLMs from the GPT, LLaMA, DeepSeek, and Qwen families, along with other reasoning models including both general domain and financial specific models. Our selection covers models of various scales, from smaller models (8B) to large-scale models (70B), and even ultra-large models like DeepSeek-R1 (671B). We also include advanced closed-source models such as GPT-4o and GPT-o1 as strong baselines. More details of the models can be found in Appendix E.
@@ 4.3 Evaluation Settings
In each dataset, LLMs are presented with a context containing financial reports, tables, equations, and symbols, along with a question. Following prior work wei2022chain , we use a generic prompt (see appendix G for more details) to instruct models to answer the question.
While answers to financial tasks are often expressed in mathematical forms, variations in representation, such as differences in percentages, rounding conventions, or numeric formatting, can lead to discrepancies in surface-level comparisons. To address this, we opt out exact match metrics and instead adopt an LLM-as-judge evaluation paradigm. Specifically, we follow the answer extraction methodology proposed by chen2023theoremqa in TheoremQA, extracting the numerical results from model outputs. The extracted answers are then compared against the ground truth to assess correctness, allowing for flexible yet rigorous evaluation of generated responses. Please see appendix D.2 for more details of the evaluation setting.
@@ 5 Results
Next, we present the results of these LLMs on FinReason and discuss our main findings through the results for the research questions listed in section  1.
Table 3: Overall performance of different LLMs on four financial reasoning datasets.
@@
Llama-4-Scout 70.45 52.00 88.89 0.67 53.00
Llama-3-70B-Instruct 58.92 41.00 56.67 13.67 42.57
@@
Dianjin-R1-7B 60.20 35.00 83.33 14.67 48.30
Fino1-8B 73.03 56.00 84.44 26.33 59.95
Fino1-14B 74.18 55.00 87.78 27.33 61.07
Table 3 presents the performance of evaluated LLMs across four financial reasoning benchmarks: FinQA, DM-Simplong, XBRL-Math, and DM-Complong. Our Fin-o1-14B achieves a score of 61.07, ranking second only to Deepseek-V3 despite having substantially fewer parameters. Notably, Fin-o1-14B surpasses larger-scale models like LLaMA-3.3-70B-Instruct, Qwen2.5-72B-Instruct, even GPT-o1-preview, GPT-o3-mini, and GPT-4.5, underscoring its superior financial reasoning capabilities. Fin-o1-8B also performs competitively, attaining an average score of 58.58, exceeding most models with 70B parameters or fewer. Both Fin-o1-8B and Fin-o1-14B excel in FinQA, DM-Simplong, and XBRL-Math, demonstrating the impact of targeted domain adaptation in financial reasoning tasks. These gains also confirm the effectiveness of our three-stage FinCoT construction and the two-phase SFT + RL training strategy.
General reasoning LLMs fail in financial reasoning.
While general reasoning models exhibit significant improvements in general mathematical reasoning tasks, they struggle to effectively transfer these gains to financial reasoning tasks, which demand specialized knowledge and reasoning over financial documents and tables. For instance, the performance of QwQ-32B and Limo declined notably compared to Qwen2.5-32B-Instruct, dropping from 56.17% to 52.92% and 46.22%, respectively. Similarly, although S1-32B demonstrated overall performance gains, its effectiveness in FinQA, Simplong and Complong decreased. These results highlight the persistent gap between general and financial reasoning due to the complexities of interpreting financial tables and structures.
Incomprehensive reasoning data limits consistent improvements. While our model demonstrates consistent improvements over the backbone model Qwen3 across all datasets, other financial models do not exhibit such stable improvements. For instance, while Fin-R1 and Dianjin-R1-7B achieved higher scores in FinQA, their performance in long-context reasoning tasks dropped significantly compared with their backbone model Qwen-2.5-7B-Instruct. On Simplong, the backbone model scored 41%, but Fin-R1 and Dianjin-R1-7B fell to 37% and 35%, respectively. A similar decline is observed on Complong, indicating the challenges in maintaining long context reasoning performance. These declines are attributable to CoT datasets used in these works, which focused on shorter contexts and knowledge-based QA rather than reasoning over contexts, limiting the models’ exposure to reasoning over extended financial contexts.
Table 4: Ablation study of Fin-o1-8B with different SFT training data and RL methods.
Models FinQA DM-Simplong XBRL-Math DM-Complong Average
Backbone model
Qwen3-8B 62.11 46.00 83.33 17.67 52.28
SFT model
Fin-o1-8B (SFT with original QA pair) 66.55 50.00 74.44 16.67 51.92
Fin-o1-8B (SFT with all CoT data) 71.32 51.00 81.11 21.67 56.28
Fin-o1-8B (SFT with filtered CoT) 71.78 57.00 82.22 23.33 58.58
SFT + RL model
Fin-o1-8B (SFT + PPO) 72.05 54.00 83.33 23.00 58.10
Fin-o1-8B (SFT + DPO) 69.74 51.00 72.22 25.00 54.49
Fin-o1-8B (SFT + GRPO) 73.03 56.00 84.44 26.33 59.95
Table 4 presents additional results evaluating the effectiveness of various RL methods in financial reasoning tasks and further verifying the effectiveness of FinCoT. Based on the results, we find that:
GRPO Outperforms Other RL Methods in Financial Reasoning.
GRPO achieves the most substantial gains, driven by its multi-faceted reward function that emphasizes correct reasoning and long-context comprehension. Notably, in DM-Complong, where other models like Fin-R1 and S1-32B falter, the SFT+GRPO model reaches 26.33% with just 8B parameters—a 9% improvement over its backbone model. In contrast, PPO shows gains in FinQA and XBRL-Math but declines in Simplong and Complong, indicating weaker long-context reasoning. For DPO, while the performance in DM-Complong improves, performance in simpler tasks declines. This may stem from the design of positive-negative sample pairs, where the negative path is curated from the last-mistake path and the positive path is generated by allowing the model to take one additional reasoning step. This approach may inadvertently encourage excessive reasoning, benefiting complex tasks but potentially causing overthinking in simpler tasks.
We further compare the impact of fine-tuning the backbone model using original QA pairs without CoT reasoning and with unfiltered CoT data. Training without CoT data slightly improves basic financial reasoning tasks like FinQA, but leads to performance drops in XBRL-Math and Complong. This decline is particularly pronounced in XBRL-Math, where the lack of step-by-step reasoning impedes the model’s ability to handle financial equations, hindering its generalization to equation-heavy tasks. In contrast, using the unfiltered FinCoT dataset demonstrates the effectiveness of exposing the model to more complex reasoning paths, underscoring the importance of reasoning-focused training for financial tasks.
@@ 6 Conclusion
In this work, we address the critical yet underexplored challenge of financial reasoning for LLMs. We present three key contributions to bridge existing gaps: (1) the development of FinCoT, the first publicly available CoT dataset specifically curated for financial reasoning tasks; (2) the introduction of Fin-o1, the first open-source financial reasoning model trained through a two-phase framework combining supervised fine-tuning and RL methods; and (3) the establishment of FinReason, a comprehensive benchmark designed to systematically evaluate financial reasoning across essential components such as multi-table analysis, long-context reasoning, and equation-based tasks.
Our extensive evaluations demonstrate that general reasoning models struggle to adapt to financial contexts, revealing significant overall performance drops. In contrast, our Fin-o1 models, trained on FinCoT with GRPO, achieve consistent gains across diverse financial reasoning tasks, highlighting the critical role of domain-specific CoT supervision and structured RL optimization. These findings underscore the need for targeted datasets, specialized models, and advanced training strategies to effectively address the complexities of financial reasoning, setting a new standard for future research in this domain.
@@ Appendix A Related Work
A.1 General-Purpose and Reasoning LLMs
Recent years have witnessed significant advancements in large language models which demonstrated remarkable capabilities across various tasks, including natural language understanding, generation, and complex reasoning. Building upon these foundations, researchers have developed specialized reasoning-enhanced models such as DeepSeek-R1 and GPT-o1, which incorporate techniques like chain-of-thought prompting [35], self-reflection [44], and reinforcement learning from human feedback [23]. These models have shown superior performance in mathematical reasoning [16], symbolic manipulation [7], and logical inference tasks [2].